Determine the expected count for each outcome.

nequals=906906

i

1

2

3

4

pi

0.160.16

0.320.32

0.190.19

0.330.33

The expected count for outcome 1 is

nothing .

(Round to two decimal places as needed.)

The expected count for outcome 2 is nothing

(Round to two decimal places as needed.)

The expected count for outcome 3 is nothing

(Round to two decimal places as needed.)

The expected count for outcome 4 is nothing

(Round to two decimal places as needed.)
Answer #1

Given that, n = 906

i) n * P(1) = 906 * 0.16 = 144.96

Therefore, The expected count for outcome 1 is 144.96

ii) n * P(2) = 906 * 0.32 = 289.92

Therefore, The expected count for outcome 2 is 289.92

iii) n * P(3) = 906 * 0.19 = 172.14

Therefore, The expected count for outcome 3 is 172.14

iv) n * P(4) = 906 * 0.33 = 298.98

Therefore, The expected count for outcome 4 is 298.98